# Benchmarking LLMs via Uncertainty Quantification

### v1

@@ Abstract
The proliferation of open-source Large Language Models (LLMs) from various institutions has highlighted the urgent need for comprehensive evaluation methods. However, current evaluation platforms, such as the widely recognized HuggingFace open LLM leaderboard, neglect a crucial aspect – uncertainty, which is vital for thoroughly assessing LLMs. To bridge this gap, we introduce a new benchmarking approach for LLMs that integrates uncertainty quantification. Our examination involves eight LLMs (LLM series) spanning five representative natural language processing tasks. Additionally, we introduce an uncertainty-aware evaluation metric, UAcc, which takes into account both prediction accuracy and prediction uncertainty. Our findings reveal that: I) LLMs with higher accuracy may exhibit lower certainty; II) Larger-scale LLMs may display greater uncertainty compared to their smaller counterparts; and III) Instruction-finetuning tends to increase the uncertainty of LLMs. By taking uncertainty into account, our new UAcc metric can either amplify or diminish the relative improvement of one LLM over another and may even change the relative ranking of two LLMs. These results underscore the significance of incorporating uncertainty in the evaluation of LLMs. [^1] Our implementation available at https://github.com/smartyfh/LLM-Uncertainty-Bench
@@ 1 Introduction
Large Language Models (LLMs) have gained significant traction within both academia and industry in recent two years, with numerous organizations and companies open-sourcing their versions of LLMs (Chang et al., 2023; Zhao et al., 2023; Kaddour et al., 2023; Yang et al., 2023a). LLMs are highly versatile, demonstrating capabilities in various tasks such as question answering, document summarization, and dialogue systems. Given the growing interest and advancements in LLMs, it is crucial to establish appropriate methods for evaluating their performance (Liang et al., 2022; Ye et al., 2023; Chang et al., 2023). However, conducting a comprehensive evaluation of LLMs remains a challenging endeavor (Guo et al., 2023).
@@
To address this challenge, several open leaderboards such as the popular HuggingFace open LLM leaderboard [^2] https://huggingface.co/spaces/HuggingFaceH4/open_llm_leaderboard, OpenCompass (Contributors, 2023), Chatbot Arena (Zheng et al., 2023), and FlagEval (BAAI, 2023) have emerged, providing a comparative analysis of LLM performance. Despite their usefulness, these leaderboards possess a significant limitation: They do not take into account the uncertainty of LLMs. For example, the HuggingFace open LLM leaderboard only utilizes accuracy as the evaluation metric. However, as demonstrated in Figure 1, two LLMs may achieve identical accuracy scores but exhibit different levels of uncertainty regarding the question. This is analogous to students taking exams of multiple-choice questions, where two students may select the same answer but actually possess distinct degrees of uncertainty or comprehension about the question. Consequently, it is necessary to incorporate uncertainty into the evaluation process to achieve a more comprehensive assessment of LLMs.
In this paper, we propose the utilization of conformal prediction (Vovk et al., 2005; Angelopoulos and Bates, 2021) as the method to quantify uncertainty in LLMs. Compared to alternative methods such as Bayesian variational inference (Hu et al., 2023), conformal prediction offers multiple advantages including ease of implementation, high efficiency, and a statistically rigorous estimation of uncertainty rather than a heuristic approximation. Therefore, conformal prediction can serve as a practical and principled means for assessing the uncertainty of LLMs.
Specifically, we benchmark eight open-source LLMs (LLM series) across five typical Natural Language Processing (NLP) tasks, namely question answering, reading comprehension, commonsense inference, dialogue response selection, and document summarization. Although some of these tasks (e.g., document summarization) are inherently generative, it is challenging to develop a deterministic and reproducible method for quantifying the uncertainty within the generated text due to randomness in the generation process. To circumvent this issue, we convert all tasks into multiple-choice questions, with the objective of each task being to select the correct option from the provided choices. Our empirical results reveal the following observations:
@@
Moreover, we introduce a novel evaluation metric, UAcc, which takes into account both prediction accuracy and prediction uncertainty. Our findings indicate that UAcc can enlarge or shrink the relative improvement of one LLM over another and may even alter the relative ranking of two LLMs. This highlights the significance of uncertainty as a crucial factor in benchmarking LLMs.
@@ 2 Related Work
Uncertainty quantification (Abdar et al., 2021; Gawlikowski et al., 2023) has been an active area of research in both machine learning and NLP due to its importance in real-world applications such as decision making, risk assessment, human-AI collaboration, and so on. Typical uncertainty quantification methods include confidence-based methods (Hu et al., 2023), Bayesian methods (Kwon et al., 2020), and ensemble methods (Rahaman et al., 2021). Confidence-based methods such as entropy can be sensitive to poor calibration and may not fully capture models’ underlying uncertainties (Minderer et al., 2021). Bayesian methods and ensemble methods usually suffer from high computational complexity (Gawlikowski et al., 2023), making them not suitable for assessing the uncertainty of LLMs.
@@
Recently, there has been a growing interest in applying conformal prediction for uncertainty quantification (Angelopoulos and Bates, 2021; Kumar et al., 2023; Quach et al., 2023; Lu et al., 2023). For example, conformal prediction has been applied to part-of-speech prediction (Dey et al., 2022), paraphrase detection (Giovannotti and Gammerman, 2021), and fact verification (Fisch et al., 2020). Similar to the process of elimination used by students during exams, conformal prediction identifies a subset of potential labels in classification tasks by excluding improbable labels, which is statistically guaranteed to contain the true label, and quantifies uncertainty as the size of this subset (Angelopoulos et al., 2020). The coverage guarantee property makes conformal prediction a highly robust uncertainty quantification method. In addition, conformal prediction is non-parametric and computationally efficient. Hence, it is a favorable choice in the context of LLMs.
@@
Evaluating the performance of LLMs is a crucial aspect of their development and deployment (Huang et al., 2023). Current studies focus on assessing LLMs from different angles using specific datasets, such as MMLU (Hendrycks et al., 2020) for knowledge, HellaSwag (Zellers et al., 2019) for reasoning, HaluEval Li et al. (2023a) for hallucination, and BOLD (Dhamala et al., 2021) for fairness. Besides, evaluation platforms like HuggingFace open LLM leaderboard and Chatbot Arena (Zheng et al., 2023) have also been developed to facilitate comparisons among LLMs. Despite these efforts, the critical aspect of uncertainty in LLMs remains underexplored. More recently, some research has begun to consider uncertainty in LLMs (Xiong et al., 2023; Yang et al., 2023b; Lin et al., 2023; Chen and Mueller, 2023; Wagle et al., 2023). However, these approaches are heuristic and lack a standardized methodology for benchmarking purposes. In contrast, our utilization of conformal prediction can provide a robust and systematic evaluation of uncertainty in LLMs.
@@ 3 Background of Conformal Prediction
Formally, let $f$ be a model that classifies an input $X$ into $K$ pre-defined classes, represented by $\mathcal{Y}=\{1,\dots,K\}$. To measure the uncertainty of $f$, for any given test instance $X_{test}$ and its corresponding true label $Y_{test}$, conformal prediction produces a prediction set of labels $\mathcal{C}(X_{test})\subset\mathcal{Y}$ such that
$p(Y_{test}\in\mathcal{C}(X_{test}))\geq 1-\alpha,$ (1)
@@
Equation (1) requires that the generated prediction set should contain the true label $Y_{test}$ with a probability no smaller than $1-\alpha$. This coverage guarantee requirement can be achieved with the aid of a small amount of held-out calibration data $\mathcal{D}_{cal}=\{(X^{(1)}_{cal},Y^{(1)}_{cal}),\dots,(X^{(n)}_{cal},Y^{(n)}_{cal})\}$, where $n$ denotes the number of data points in the calibration set [^3] It is also required that the test data points are drawn from the same distribution as the calibration data.. More specifically, conformal prediction works in the following process (Angelopoulos and Bates, 2021) to create the prediction set:
@@
    Compute conformal scores on the calibration set $s_{1}=s(X^{(1)}_{cal},Y^{(1)}_{cal}),\dots,s_{n}=(X^{(n)}_{cal},Y^{(n)}_{cal})$ and calculate a threshold $\hat{q}$ as the $\frac{\lceil(n+1)(1-\alpha)\rceil}{n}$ quantile of the calibration scores,
@@
    Construct the prediction set for each test instance $X_{test}$ as
    $\mathcal{C}(X_{test})=\{Y\in\mathcal{Y}:s(X_{test},Y)\leq\hat{q}\}.$ (3)
@@
$s(X,Y)=\sum_{\{j\in\mathcal{Y}:f(x)_{j}\geq f(x)_{Y}\}}f(X)_{j},$ (5)
where $f(X)_{j}$ represents the softmax score corresponding to the label $j\in\mathcal{Y}$. Equation (5) is equivalent to summing the ranked scores of each label, from the higher to the lower, until reaching the true label. Compared to LAC, APS leverages the softmax scores of all labels, not just the true label. It addresses the limitation of LAC but suffers from, on average, larger prediction sets.
Figure 2: The overall process of applying conformal prediction for uncertainty quantification in LLMs. (a) Five distinct tasks are considered, and a dataset comprising 10,000 instances is prepared for each task. (b) Each data instance is transformed into a multiple-choice question, and eight LLMs (LLM series) are prompted to generate predicted probabilities for the given options. (c) Each dataset is divided into a calibration set and a test set, followed by the application of conformal prediction to generate prediction sets for test set instances. For illustrative purposes, demonstrations in the prompt input are excluded, and solely the process of constructing prediction sets utilizing the LAC conformal score function is demonstrated. In addition, only four options of the question are presented.
@@ 4 Evaluation Tasks and Datasets
LLMs have demonstrated remarkable capabilities across various aspects (Guo et al., 2023; Naveed et al., 2023). It is essential to develop multiple tasks to evaluate their performance comprehensively. For this purpose, we consider five typical NLP tasks, including question answering, reading comprehension, commonsense inference, dialogue response selection, and document summarization. For each task, we prepare a dataset with 10,000 instances. In addition, we formulate each task as a multiple-choice question answering (MCQA) task and the objective is to select the only correct answer out of six possible options (i.e. A, B, C, D, E, and F).
@@ 5.1 Prompting Strategies
Following previous works (Hendrycks et al., 2020; Gao et al., 2023; Zhang et al., 2023; Chen et al., 2023; Zheng et al., 2023), we rely on prompt engineering rather than supervised finetuning as the testing approach to evaluating the performance of LLMs on each task. However, our preliminary results show that LLMs are sensitive to prompts. In this regard, we consider three prompting strategies, including Base Prompt, Shared Instruction Prompt, and Task-specific Instruction Prompt, to reduce the influence of LLMs’ sensitivity to different prompts, thereby ensuring a fairer comparison.
@@ 5.2 Evaluation Metrics
We evaluate LLMs from two perspectives, namely prediction accuracy and prediction uncertainty. For prediction accuracy, we adopt the commonly used metric of Accuracy (Acc), which measures the proportion of test instances whose true label has the highest softmax score. To evaluate prediction uncertainty, we use Set Size (SS), which calculates the average length of prediction sets of all test instances. These prediction sets are obtained using conformal prediction, as explained in Section 3.
In addition to Acc and SS, we propose a new metric, Uncertainty-aware Accuracy (UAcc), which takes into account both the prediction accuracy and uncertainty. UAcc is calculated as follows:
$UAcc=\frac{Acc}{SS}\sqrt{|\mathcal{Y}|},$ (6)
where the incorporation of $\sqrt{|\mathcal{Y}|}$ allows UAcc to be smaller or larger than Acc, depending on the level of uncertainty in the predictions.
@@ 6.1 Setup
In our experiments, we set the error rate $\alpha$ to 0.1, implying that the prediction set should include the true label with a probability of at least 0.9. In order to better excite the ability of LLMs, we incorporate examples or demonstrations in the prompt, adhering to the in-context learning paradigm (Dong et al., 2022). Specifically, we provide five demonstrations for QA, RC, and CI tasks. For the DRS task, we utilize three demonstrations, while we use only one demonstration for the DS task due to constraints on input length and inference cost. The maximum input length for all tasks is set to 2048 tokens. In addition, for each task, we allocate 50% of the data as the calibration set and the remaining 50% as the test set. We report results on the test set. These results represent the average value obtained from the two conformal score functions, namely, LAC and APS, as well as the three prompting strategies. It is noteworthy that while both LAC and APS fulfill the coverage guarantee requirement, they may produce prediction sets of varying sizes. By taking the average value of these score functions, we aim to mitigate the influence of different score functions on the evaluation of uncertainty, thereby ensuring a more rigorous and reliable assessment.
LLMs Coverage Rate (%) Acc (%) $\uparrow$
@@
Qwen-14B 92.58 95.20 95.29 91.92 89.71 92.94 64.25(1) 91.52(1) 91.00(1) 73.90(1) 49.33(4) 74.00(1)
Yi-6B 91.30 94.06 92.35 91.36 91.38 92.09 57.57(3) 85.99(2) 76.50(2) 58.72(3) 66.06(1) 68.97(2)
Mistral-7B 93.00 92.91 89.94 91.02 92.05 91.78 60.44(2) 81.94(4) 62.93(4) 53.21(4) 62.16(2) 64.14(3)
Llama-2-13B 92.59 93.15 90.50 90.92 90.82 91.60 52.52(5) 77.23(5) 59.66(5) 52.65(5) 60.05(3) 60.42(4)
Qwen-7B 92.79 94.02 91.53 92.43 89.56 92.07 55.21(4) 83.89(3) 63.70(3) 64.04(2) 32.53(8) 59.87(5)
InternLM-7B 90.68 93.28 90.10 90.40 90.34 90.96 48.37(6) 73.86(6) 46.21(6) 43.72(6) 34.38(7) 49.31(6)
Llama-2-7B 91.37 90.69 90.97 89.60 90.04 90.53 45.60(8) 65.79(7) 43.05(7) 32.61(8) 45.60(5) 46.53(7)
DeepSeek-7B 91.18 89.95 90.16 90.89 90.21 90.48 45.65(7) 65.39(8) 42.66(8) 33.50(7) 42.15(6) 45.87(8)
MPT-7B 89.79 90.54 90.12 90.80 89.71 90.19 29.49(9) 31.69(9) 25.50(9) 24.38(10) 24.86(9) 27.18(9)
Falcon-7B 90.04 89.95 89.82 90.46 90.71 90.19 23.75(10) 24.98(10) 24.91(10) 25.86(9) 24.69(10) 24.84(10)
LLMs SS $\downarrow$ UAcc (%) $\uparrow$
QA RC CI DRS DS Avg. QA RC CI DRS DS Avg.
Qwen-14B 2.80(1) 1.74(1) 2.02(2) 1.94(1) 2.37(3) 2.17(1) 57.83(1) 157.52(1) 147.13(1) 97.70(1) 51.22(4) 102.28(1)
Yi-6B 3.20(4) 1.92(3) 1.88(1) 2.85(5) 1.96(1) 2.36(2) 45.18(3) 132.61(2) 103.41(2) 50.97(3) 85.49(1) 83.53(2)
Mistral-7B 2.80(1) 1.75(2) 2.48(4) 2.71(4) 2.40(4) 2.43(3) 54.60(2) 124.71(3) 62.45(4) 48.18(5) 64.25(3) 70.84(3)
Llama-2-13B 3.06(3) 2.24(6) 2.72(5) 2.55(3) 2.24(2) 2.56(4) 42.53(4) 92.46(5) 53.82(5) 50.52(4) 66.02(2) 61.07(5)
Qwen-7B 3.26(6) 2.15(4) 2.28(3) 2.51(2) 2.92(5) 2.63(5) 42.45(5) 118.10(4) 69.47(3) 64.42(2) 27.28(7) 64.34(4)
InternLM-7B 3.49(8) 2.19(5) 3.28(8) 3.63(9) 4.47(10) 3.41(8) 34.17(7) 86.84(6) 34.56(6) 29.73(6) 18.87(8) 40.83(6)
Llama-2-7B 3.20(4) 2.39(7) 3.27(7) 3.26(6) 3.30(7) 3.09(6) 34.97(6) 67.92(7) 32.25(8) 24.50(7) 33.91(5) 38.71(7)
DeepSeek-7B 3.34(7) 2.77(8) 3.06(6) 3.40(7) 3.08(6) 3.13(7) 33.63(8) 58.50(8) 34.23(7) 24.11(8) 33.52(6) 36.80(8)
MPT-7B 3.53(9) 3.46(9) 3.60(9) 3.59(8) 3.66(8) 3.57(9) 20.44(9) 22.43(9) 17.36(9) 16.66(10) 16.63(9) 18.70(9)
Falcon-7B 3.90(10) 3.60(10) 3.66(10) 3.64(10) 3.92(9) 3.75(10) 14.90(10) 17.01(10) 16.66(10) 17.41(9) 15.42(10) 16.28(10)
Table 1: Results of LLMs with sizes ranging from 6B to 14B. These results represent the mean values of LAC and APS. The "Avg." column denotes the average performance across the five tasks. The small number in parentheses next to each score indicates the ranking of the LLM for that specific task. The relative ranking of LLMs is also visually demonstrated by the depth of color, with darker colors signifying higher rankings. Note that UAcc takes values in the range $[0,\sqrt{|\mathcal{Y}|}]$, thus can be larger than 1.

### v2

@@ Abstract
The proliferation of open-source Large Language Models (LLMs) from various institutions has highlighted the urgent need for comprehensive evaluation methods. However, current evaluation platforms, such as the widely recognized HuggingFace open LLM leaderboard, neglect a crucial aspect – uncertainty, which is vital for thoroughly assessing LLMs. To bridge this gap, we introduce a new benchmarking approach for LLMs that integrates uncertainty quantification. Our examination involves nine LLMs (LLM series) spanning five representative natural language processing tasks. Our findings reveal that: I) LLMs with higher accuracy may exhibit lower certainty; II) Larger-scale LLMs may display greater uncertainty compared to their smaller counterparts; and III) Instruction-finetuning tends to increase the uncertainty of LLMs. These results underscore the significance of incorporating uncertainty in the evaluation of LLMs. [^1] Our implementation available at https://github.com/smartyfh/LLM-Uncertainty-Bench
@@ 1 Introduction
Large Language Models (LLMs) have gained significant traction within both academia and industry, with numerous organizations and companies open-sourcing their versions of LLMs (Chang et al., 2023; Zhao et al., 2023; Kaddour et al., 2023; Yang et al., 2023a). LLMs are highly versatile, demonstrating capabilities in various tasks such as question answering, document summarization, dialogue systems, and machine translation (Ye et al., 2023a; Pang et al., 2024). Given the growing interest and advancements in LLMs, it is crucial to establish appropriate methods for evaluating their performance (Liang et al., 2022; Ye et al., 2023b; Chang et al., 2023). However, conducting a comprehensive evaluation of LLMs remains a challenging endeavor (Guo et al., 2023; Zheng et al., 2024).
@@
To address this challenge, several open leaderboards such as the popular HuggingFace open LLM leaderboard [^2] https://huggingface.co/spaces/HuggingFaceH4/open_llm_leaderboard, OpenCompass (Contributors, 2023), Chatbot Arena (Zheng et al., 2024), and FlagEval (BAAI, 2023) have emerged, providing a comparative analysis of LLM performance. Despite their usefulness, these leaderboards possess a significant limitation: They do not take into account the uncertainty of LLMs. For example, the HuggingFace open LLM leaderboard only utilizes accuracy as the evaluation metric. However, as demonstrated in Figure 1, two LLMs may achieve identical accuracy scores but exhibit different levels of uncertainty regarding the question. This is analogous to students taking exams of multiple-choice questions, where two students may select the same answer but actually possess distinct degrees of uncertainty or comprehension about the question. Consequently, it is necessary to incorporate uncertainty into the evaluation process to achieve a more comprehensive assessment of LLMs.
In this paper, we propose the utilization of conformal prediction (Vovk et al., 2005; Angelopoulos and Bates, 2021) as the method to quantify uncertainty in LLMs. Compared to alternative methods such as Bayesian variational inference (Hu et al., 2023), conformal prediction offers multiple advantages including ease of implementation, high efficiency, distribution-free and model-agnostic, and a statistically rigorous estimation of uncertainty rather than a heuristic approximation. Hence, conformal prediction can serve as a practical and principled means for assessing the uncertainty of LLMs.
Specifically, we benchmark nine open-source LLMs (LLM series) across five typical Natural Language Processing (NLP) tasks, namely question answering, reading comprehension, commonsense inference, dialogue response selection, and document summarization. Given that most existing open leaderboards and benchmarking datasets (Liu et al., 2024) focus on multiple-choice tasks, we also adopt the multiple-choice question setting for all tasks [^3] We leave the benchmarking of LLM uncertainty in the free-form text generation setting as our future work.. Although some of these tasks (e.g., document summarization) are inherently generative, it is challenging to develop a deterministic and reproducible method for quantifying the uncertainty within the generated text due to randomness in the generation process. Instead, we convert all tasks into multiple-choice questions, with the objective of each task being to select the correct option from the provided choices. Our empirical results reveal the following observations:
@@ 2 Related Work
Uncertainty quantification (Fomicheva et al., 2020; Abdar et al., 2021; Gawlikowski et al., 2023) has been an active area of research in both machine learning and NLP due to its importance in real-world applications such as decision making, risk assessment, human-AI collaboration, and so on. Typical uncertainty quantification methods include confidence-based methods (Hu et al., 2023), Bayesian methods (Kwon et al., 2020), and ensemble methods (Rahaman et al., 2021). Confidence-based methods such as entropy can be sensitive to poor calibration and may not fully capture models’ underlying uncertainties (Minderer et al., 2021). Bayesian methods and ensemble methods usually suffer from high computational complexity (Gawlikowski et al., 2023), making them not suitable for assessing the uncertainty of LLMs.
@@
Recently, there has been a growing interest in applying conformal prediction for uncertainty quantification (Angelopoulos and Bates, 2021; Kumar et al., 2023; Quach et al., 2023; Lu et al., 2023). For example, conformal prediction has been applied to part-of-speech prediction (Dey et al., 2022), paraphrase detection (Giovannotti and Gammerman, 2021), and fact verification (Fisch et al., 2020). Similar to the process of elimination used by students during exams, conformal prediction identifies a subset of potential labels in classification tasks by excluding improbable labels, which is statistically guaranteed to contain the true label, and quantifies uncertainty as the size of this subset (Angelopoulos et al., 2020). The coverage guarantee property makes conformal prediction a highly robust uncertainty quantification method. In addition, conformal prediction is non-parametric, distribution-free, model-agnostic, and computationally efficient. Therefore, it is a favorable choice in the context of LLMs.
@@
Evaluating the performance of LLMs is a crucial aspect of their development and deployment (Huang et al., 2023). Current studies assess LLMs from different angles using specific datasets, such as MMLU (Hendrycks et al., 2020) for knowledge, HellaSwag (Zellers et al., 2019) for reasoning, HaluEval Li et al. (2023a) for hallucination, GSM8K (Cobbe et al., 2021) for math, and BOLD (Dhamala et al., 2021) for fairness. Besides, evaluation platforms like HuggingFace open LLM leaderboard and Chatbot Arena (Zheng et al., 2024) have also been developed to facilitate comparisons among LLMs. Despite these efforts, the critical aspect of uncertainty in LLMs remains underexplored. More recently, some research has begun to consider uncertainty in LLMs (Xiong et al., 2023; Yang et al., 2023b; Lin et al., 2023; Chen and Mueller, 2023; Wagle et al., 2023; Fadeeva et al., 2023). However, these approaches such as the sampling-based semantic entropy (Kuhn et al., 2023) are heuristic and lack a standardized methodology for benchmarking purposes. In contrast, our utilization of conformal prediction can provide a robust and systematic evaluation of uncertainty.
@@ 3 Background of Conformal Prediction
Formally, let $f$ be a model that classifies an input $X$ into $K$ pre-defined classes, represented by $\mathcal{Y}=\{1,\dots,K\}$. To measure the uncertainty of $f$, for any given test instance $X_{t}$ and its corresponding true label $Y_{t}$, conformal prediction produces a prediction set of labels $\mathcal{C}(X_{t})\subset\mathcal{Y}$ such that
$p(Y_{t}\in\mathcal{C}(X_{t}))\geq 1-\alpha,$ (1)
@@
Equation (1) requires that the generated prediction set should contain the true label $Y_{t}$ with a probability no smaller than $1-\alpha$. This coverage guarantee requirement can be achieved with the aid of a small amount of held-out calibration data $\mathcal{D}_{cal}=\{(X^{(1)}_{c},Y^{(1)}_{c}),\dots,(X^{(n)}_{c},Y^{(n)}_{c})\}$, where $n$ denotes the number of data points in the calibration set [^4] It is also required that the test data points are drawn from the same distribution as the calibration data.. More specifically, conformal prediction works in the following process (Angelopoulos and Bates, 2021) to create the prediction set:
@@
    Compute conformal scores on the calibration set $s_{1}=s(X^{(1)}_{c},Y^{(1)}_{c}),\dots,s_{n}=(X^{(n)}_{c},Y^{(n)}_{c})$ and calculate a threshold $\hat{q}$ as the $\frac{\lceil(n+1)(1-\alpha)\rceil}{n}$ quantile of the calibration scores,
@@
    Construct the prediction set for each test instance $X_{t}$ as
    $\mathcal{C}(X_{t})=\{Y^{\prime}\in\mathcal{Y}:s(X_{t},Y^{\prime})\leq\hat{q}\}.$ (3)
@@
$s(X,Y)=\sum_{\{Y^{\prime}\in\mathcal{Y}:f(x)_{Y^{\prime}}\geq f(x)_{Y}\}}f(X)_{Y^{\prime}},$ (5)
where $f(X)_{Y^{\prime}}$ represents the softmax score corresponding to the label $Y^{\prime}\in\mathcal{Y}$. Equation (5) is equivalent to summing the ranked scores of each label, from the higher to the lower, until reaching the true label. Compared to LAC, APS leverages the softmax scores of all labels, not just the true label. It addresses the limitation of LAC but suffers from, on average, larger prediction sets.
Figure 2: The overall process of applying conformal prediction for uncertainty quantification in LLMs. (a) Five distinct tasks are considered, and a dataset comprising 10,000 instances is prepared for each task. (b) Each data instance is transformed into a multiple-choice question, and nine LLMs (LLM series) are prompted to generate predicted probabilities for the given options. (c) Each dataset is divided into a calibration set and a test set, followed by the application of conformal prediction to generate prediction sets for test set instances. For illustrative purposes, demonstrations in the prompt input are excluded, and solely the process of constructing prediction sets utilizing the LAC conformal score function is demonstrated. In addition, only four options of the question are presented.
@@ 4 Evaluation Tasks and Datasets
LLMs have demonstrated remarkable capabilities across various aspects (Guo et al., 2023; Naveed et al., 2023). It is essential to develop multiple tasks to evaluate their performance comprehensively. For this purpose, we consider five typical NLP tasks, including question answering, reading comprehension, commonsense inference, dialogue response selection, and document summarization. For each task, we prepare a dataset with 10,000 instances. In addition, we formulate each task as a multiple-choice question answering (MCQA) task and the objective is to select the only correct answer out of six possible options (i.e. A, B, C, D, E, and F). It is worth emphasizing that the prevailing benchmarking open leaderboards and datasets also focuses on MCQA tasks (Guo et al., 2023; Liu et al., 2024).
@@ 5.1 Prompting Strategies
Following previous works (Hendrycks et al., 2020; Gao et al., 2023; Zhang et al., 2023; Chen et al., 2023; Zheng et al., 2024), we rely on prompt engineering rather than supervised finetuning as the testing approach to evaluating the performance of LLMs on each task. However, our preliminary results show that LLMs are sensitive to prompts. In this regard, we consider three prompting strategies, including Base Prompt, Shared Instruction Prompt, and Task-specific Instruction Prompt, to reduce the influence of LLMs’ sensitivity to different prompts, thereby ensuring a fairer comparison.
@@ 5.2 Evaluation Metrics
We evaluate LLMs from two perspectives, namely prediction accuracy and prediction uncertainty. For prediction accuracy, we adopt the commonly used metric – Accuracy (Acc). To evaluate prediction uncertainty, we use Set Size (SS), which is a primary metric for conformal prediction (Angelopoulos and Bates, 2021). Let $Y_{p}$ be the prediction for the test instance $(X_{t},Y_{t})\in\mathcal{D}_{test}$. These two metrics can be calculated as follows:
$Acc=\frac{1}{|\mathcal{D}_{test}|}\sum_{(X_{t},Y_{t})\in\mathcal{D}_{test}}\mathbbm{1}(Y_{p}=Y_{t}),$ (6)
$SS=\frac{1}{|\mathcal{D}_{test}|}\sum_{(X_{t},Y_{t})\in\mathcal{D}_{test}}|\mathcal{C}(X_{t})|,$ (7)
where $\mathbbm{1}(\cdot)$ is the indicator function.
In addition to Acc and SS, we report the coverage rate (CR) to verify if the coverage guarantee requirement shown in Eq. (1) has been satisfied. The CR metric is calculated as
$CR=\frac{1}{|\mathcal{D}_{test}|}\sum_{(X_{t},Y_{t})\in\mathcal{D}_{test}}\mathbbm{1}(Y_{t}\in\mathcal{C}(X_{t})).$ (8)
@@ 6.1 Setup
In our experiments, we set the error rate $\alpha$ to 0.1, implying that the prediction set should include the true label with a probability of at least 0.9. In order to better excite the ability of LLMs, we incorporate examples or demonstrations in the prompt, adhering to the in-context learning paradigm (Dong et al., 2022). Specifically, we provide five demonstrations for QA, RC, and CI tasks. For the DRS task, we utilize three demonstrations, while we use only one demonstration for the DS task due to constraints on input length and inference cost. The maximum input length for all tasks is set to 2048 tokens. In addition, for each task, we allocate 50% of the data as the calibration set and the remaining 50% as the test set. We report results on the test set. These results represent the average value obtained from the two conformal score functions, namely, LAC and APS, as well as the three prompting strategies. It is noteworthy that while both LAC and APS fulfill the coverage guarantee requirement, they may produce prediction sets of varying sizes. By taking the average value of these two score functions, we aim to mitigate the influence of different score functions on the evaluation of uncertainty, thereby ensuring a more rigorous and reliable assessment.
LLMs QA RC CI DRS DS Avg.
Qwen-14B 92.58 95.20 95.29 91.92 89.71 92.94
Yi-6B 91.30 94.06 92.35 91.36 91.38 92.09
Gemma-7B 93.57 94.16 92.13 91.59 90.70 92.43
Mistral-7B 93.00 92.91 89.94 91.02 92.05 91.78
Llama-2-13B 92.59 93.15 90.50 90.92 90.82 91.60
Qwen-7B 92.79 94.02 91.53 92.43 89.56 92.07
InternLM-7B 90.68 93.28 90.10 90.40 90.34 90.96
Llama-2-7B 91.37 90.69 90.97 89.60 90.04 90.53
DeepSeek-7B 91.18 89.95 90.16 90.89 90.21 90.48
MPT-7B 89.79 90.54 90.12 90.80 89.71 90.19
Falcon-7B 90.04 89.95 89.82 90.46 90.71 90.19
Table 1: The CR (%) results of LLMs with sizes ranging from 6B to 14B. These results represent the mean values of LAC and APS. The "Avg." column denotes the average performance across the five tasks.
LLMs Acc (%) $\uparrow$ SS $\downarrow$
@@
Qwen-14B 64.25(1) 91.52(1) 91.00(1) 73.90(1) 49.33(4) 74.00(1) 2.80(2) 1.74(1) 2.02(2) 1.94(1) 2.37(3) 2.17(1)
Yi-6B 57.57(4) 85.99(2) 76.50(2) 58.72(4) 66.06(1) 68.97(2) 3.20(5) 1.92(4) 1.88(1) 2.85(6) 1.96(1) 2.36(2)
Gemma-7B 62.24(2) 85.29(3) 73.58(3) 66.79(2) 40.80(7) 65.74(3) 2.72(1) 1.88(3) 2.04(3) 2.14(2) 3.11(7) 2.38(3)
Mistral-7B 60.44(3) 81.94(5) 62.93(5) 53.21(5) 62.16(2) 64.14(4) 2.80(2) 1.75(2) 2.48(5) 2.71(5) 2.40(4) 2.43(4)
Llama-2-13B 52.52(6) 77.23(6) 59.66(6) 52.65(6) 60.05(3) 60.42(5) 3.06(4) 2.24(7) 2.72(6) 2.55(4) 2.24(2) 2.56(5)
Qwen-7B 55.21(5) 83.89(4) 63.70(4) 64.04(3) 32.53(9) 59.87(6) 3.26(7) 2.15(5) 2.28(4) 2.51(3) 2.92(5) 2.63(6)
InternLM-7B 48.37(7) 73.86(7) 46.21(7) 43.72(7) 34.38(8) 49.31(7) 3.49(9) 2.19(6) 3.28(9) 3.63(10) 4.47(11) 3.41(9)
Llama-2-7B 45.60(9) 65.79(8) 43.05(8) 32.61(9) 45.60(5) 46.53(8) 3.20(5) 2.39(8) 3.27(8) 3.26(7) 3.30(8) 3.09(7)
DeepSeek-7B 45.65(8) 65.39(9) 42.66(9) 33.50(8) 42.15(6) 45.87(9) 3.34(8) 2.77(9) 3.06(7) 3.40(8) 3.08(6) 3.13(8)
MPT-7B 29.49(10) 31.69(10) 25.50(10) 24.38(11) 24.86(10) 27.18(10) 3.53(10) 3.46(10) 3.60(10) 3.59(9) 3.66(9) 3.57(10)
Falcon-7B 23.75(11) 24.98(11) 24.91(11) 25.86(10) 24.69(11) 24.84(11) 3.90(11) 3.60(11) 3.66(11) 3.64(11) 3.92(10) 3.75(11)
Table 2: The Acc and SS results of LLMs with sizes ranging from 6B to 14B. These results represent the mean values of LAC and APS. The "Avg." column denotes the average performance across the five tasks. The small number in parentheses next to each score indicates the ranking of the LLM for that specific task. The relative ranking of LLMs is also visually demonstrated by the depth of color, with darker colors signifying higher rankings.